OpenStudy (anonymous):

Write an equation in standard form for a line passing through A(-2, 3) and B(3, 4)

OpenStudy (uri):

First find slope.. Slope formula: \[\frac{ (y2-y1) }{ (x2-x1) }\]

OpenStudy (anonymous):

1/5

OpenStudy (uri):

4-3/3-(-2) 1/5 correct

OpenStudy (uri):

Now do you know the formula of how to find equation?

OpenStudy (anonymous):

no

OpenStudy (uri):

gtg @Luigi0210 please help her

OpenStudy (anonymous):

so b = 17/5

OpenStudy (luigi0210):

To find an equation you have to use the point-slope formula: y-y1=m(x-x1)

OpenStudy (luigi0210):

Uri was right about finding the slope. m=slope Now you have to choose 1 point out of those two, which one would you like to use?

OpenStudy (anonymous):

(-2, 3)

OpenStudy (uri):

Oh sorryyy Formula is y-y1=m(x-x1) y1=3 x1=-2 Plug in the formula @ejune420

OpenStudy (uri):

values*

OpenStudy (anonymous):

y-3=1/5(x+2)

OpenStudy (luigi0210):

Yup, now just distribute and finish the problem :)

OpenStudy (uri):

x-2* not +

OpenStudy (luigi0210):

No she's right

OpenStudy (luigi0210):

It's (x-(-2))

OpenStudy (luigi0210):

So finish it off and show me what you get as a final answer.

OpenStudy (anonymous):

I got to y-3=1/5x+2/5 and idk what to do next

OpenStudy (uri):

Oh okay..

OpenStudy (luigi0210):

Oh uri :3 Now what you do is add 3 to both sides.

OpenStudy (anonymous):

y=1/5x+17/5

OpenStudy (luigi0210):

Yea, and there's your answer :)

OpenStudy (anonymous):

thats not an answer choice :/

OpenStudy (luigi0210):

What are the choices?

OpenStudy (anonymous):

5x-y=17 x-y=1 x-5y=19 x-5y=-17

OpenStudy (luigi0210):

Oh, standard form.. oops, sorry.

OpenStudy (luigi0210):

Okay, what we do is get y and x all on one side. SO move the x over.

OpenStudy (anonymous):

y-1/5x=17/5

OpenStudy (luigi0210):

Yes, but make sure x is the leading value

OpenStudy (anonymous):

sooo -1/5x+y=17/5

OpenStudy (luigi0210):

So it's \[-\frac{1}{5}x+y=\frac{17}{5}\]

OpenStudy (anonymous):

that's still not an answer choice though :(

OpenStudy (luigi0210):

We're not done yet, be patient :P

OpenStudy (anonymous):

hahah oh sorry :p

OpenStudy (luigi0210):

Okay, so in standard form: Ax+By=C x can not be a fraction nor negative, so we need to get rid of that -1/5, do you know how?

OpenStudy (anonymous):

ummmmm multiply by 5?

OpenStudy (luigi0210):

*-5 And yes, just do that and you'll have your solution :)

OpenStudy (anonymous):

x-5y=-17?

OpenStudy (luigi0210):

Yea :)

OpenStudy (anonymous):

thanks so much!

OpenStudy (luigi0210):

Anytime ☺